### v2

# GTransPDM: A Graph-embedded Transformer with Positional Decoupling for Pedestrian Crossing Intention Prediction

Abstract
Understanding and predicting pedestrian crossing behavioral intention is crucial for the driving safety of autonomous vehicles. Nonetheless, challenges emerge when using promising images or environmental context masks to extract various factors for time-series network modeling, causing pre-processing errors or a loss of efficiency. Typically, pedestrian positions captured by onboard cameras are often distorted and do not accurately reflect their actual movements. To address these issues, GTransPDM – a Graph-embedded Transformer with a Position Decoupling Module – was developed for pedestrian crossing intention prediction by leveraging multi-modal features. First, a positional decoupling module was proposed to decompose pedestrian lateral motion and encode depth cues in the image view. Then, a graph-embedded Transformer was designed to capture the spatio-temporal dynamics of human pose skeletons, integrating essential factors such as position, skeleton, and ego-vehicle motion. Experimental results indicate that the proposed method achieves 92% accuracy on the PIE dataset and 87% accuracy on the JAAD dataset, with a processing speed of 0.05ms. It outperforms the state-of-the-art in comparison.
Index Terms:
Pedestrian behavior, Intention prediction, Time-series model, Traffic safety.
I Introduction
Pedestrians, as vulnerable road users, are a crucial concern for the safe operation of autonomous vehicles (AVs). Accurately understanding and predicting pedestrian behavioral intentions is essential for AVs to make proactive safety decisions [1, 2]. Pedestrian crossing intention prediction (PCIP) continues to be a challenge in time series classification, where the input comprises a sequence of features derived from sensor data, and the output indicates the probability that a person will cross the road after a given prediction horizon [3] (Fig. 1).
Figure 1: Problem Definition of Pedestrian Crossing Intention Prediction (PCIP): Given $T$ frames of observations, PCIP predicts whether a pedestrian will cross the road within the next 1-2 seconds, allowing the ego-vehicle to react in advance. In this work, we developed a multi-modal fusion approach, incorporating factors such as pedestrian position, human skeleton pose, and ego-vehicle motion.
In existing research, RNN-based approaches have been employed to develop time series models, either independently or in combination with graph neural networks (GNNs), to represent skeletal information [4, 5, 6, 7, 8]. However, the role of skeleton information within the feature fusion paradigm remains controversial in current studies, underscoring the necessity for precise extraction of pose dynamics [9, 10]. To gain a deeper insight into pedestrian crossing behavior, various factors have been explored in existing approaches, showcasing improved performance by incorporating environmental context [11, 12, 13, 14]. However, these methods may lead to errors in image processing or instance mask inference, and may also become computationally intensive. Pedestrian trajectories provide valuable insights into pedestrian behavior and have been extensively utilized in PCIP [15, 16, 17, 18]. Yet, since pedestrian positions in the image are relative to the ego-vehicle, this can hinder a model’s ability to accurately identify pedestrian movements in real-world scenarios. Typically, a pedestrian’s lateral shift toward the ego-vehicle is a key indicator of a potential crossing. However, the position may be distorted due to the motion of the ego-vehicle, making it challenging to pinpoint their actual movements. Moreover, the distance between pedestrians and the ego-vehicle greatly influences the crossing decision [19, 20]. This factor, though critical, has often been neglected in current studies due to the complexity of depth estimation.
To address the above issues, a Graph-embedded Transformer with a Position Decoupling Module (GTransPDM) was developed by merging multi-modal features. Based on the observation that crossing pedestrians in conflict with the ego-vehicle inevitably exhibit lateral movement toward it, we propose a simple yet intuitive PDM to enhance pedestrian lateral movement recognition using reference lines. Then, the area ratio of the bounding boxes was proposed as a proxy representation of depth variation, demonstrating strong consistency with actual depth measurements. Finally, residual graph convolutional blocks with learnable edge importance were designed within a Transformer-based architecture for PCIP, enabling the consideration of spatio-temporal variations in human pose skeletons. The main contributions are summarized as follows.
  • We are the first to address the motion disparity between ego-vehicles and pedestrians to improve pedestrian position representations from on-board sensors. Building on this insight, we introduce PDM, a method that decouples the movements of AVs and pedestrians to enhance the recognition of pedestrians approaching the vehicle. Moreover, we leverage the area ratio of bounding box as depth cues instead of costly depth estimation.
  • We propose a fast Transformer-based multi-modal fusion framework to enhance PCIP by integrating data from various sensors. Notably, we develop residual graph convolutional blocks with learnable edge importance to capture the spatio-temporal dynamics of the human pose skeleton, which allows for accurate modeling of joint interactions and motion patterns.
  • Extensive experiments on the large-scale PIE and JAAD datasets demonstrate the superior performance of our proposed method, achieving an inference time of 0.05 ms using a single GPU, outperforming existing methods.
II Methodology
Figure 2: Overview of the proposed framework.
II-A Overview
An overview of the proposed framework is presented in Fig. 2. In the presented framework, the encoding of the network was performed using position, pose, and ego-vehicle motion toward pedestrian $i$ across $T$ observation frames. Each encoder was referred to as $\mathbf{X}_{pe}$, $\mathbf{X}_{ke}$, and $\mathbf{X}_{ev}$, respectively. In positional encoder, PDM $P_{i}$ was developed to decompose true lateral movements of the pedestrian and reflect depth changes, which were then integrated with displacement $D_{i}$ and instantaneous velocity $V_{i}$ to further represent the positional shift in image view. Then, the ego-vehicle’s motion, which indicates potential conflict with a pedestrian, was characterized by its speed $S_{i}$ and acceleration $Acc_{i}$. In addition, several Graph Convolutional Network (GCN) blocks equipped with residual connections were layered to capture spatial dynamics of the pedestrian pose $K_{i}$. Finally, these encoded representations were combined and fed into a Transformer encoder for temporal modeling, followed by an MLP head for classification. As described above, the PCIP problem in this study can be outlined as follows.
$\hat{Y}_{i}=\operatorname{argmax}\mathcal{P}\left(\mathbf{X}_{pe}\left[P_{i},D_{i},V_{i}\right],\mathbf{X}_{ke}\left[K_{i}\right],\mathbf{X}_{ev}\left[S_{i},Acc_{i}\right]\right)$ (1)
II-B Network Architecture
Positional Encoder. In this study, the positional encoder includes displacement $D_{i}=\left\{x_{t}-x_{0},y_{t}-y_{0}\right\}\in\mathbb{R}^{T\times 2}$, instantaneous velocity $V_{i}=\left\{x_{t}-x_{t-1},y_{t}-y_{t-1}\right\}\in\mathbb{R}^{T\times 2}$, and $P_{i}$ from the proposed PDM, where $x_{t}$ and $y_{t}$ are the center coordinates of pedestrian bounding boxes, and $t\in(0,...,T)$. $D_{i}$ and $V_{i}$ collectively describe the overall and frame-by-frame displacement of the pedestrian relative to the ego-vehicle. Our proposed PDM is shown in Fig. 2 (a). Firstly, the region of interest (ROI) for the crossing behavior was enclosed using two reference lines $(x_{l},y_{l})$. Then, the disparity between the pedestrian center and the reference line at frame $t$ was calculated, namely $(\Delta x_{t},\Delta y_{t})$, where $\Delta x_{t}=x_{t}-x_{l}$, $\Delta y_{t}=y_{t}-y_{l}$ to indicate position disparity between them. Finally, the area ratio $R$ was employed to combine with the position variations in PDM to capture the depth shift in image view. The whole features in $P_{i}$ can be represented as:
$\begin{gathered}P_{i}=\left\{\Delta x_{t}-\Delta x_{t-1},\Delta y_{t}-\Delta y_{t-1},R\right\}\in\mathbb{R}^{T\times 3}\\ R=\left(\frac{A_{t}}{A_{t-1}}-1\right)\times\alpha\end{gathered}$ (2)
where: $A_{t}$ is the bounding box area at frame $t$. $\alpha$ is a scaling factor and was set as $\alpha=100$. Each feature in $\mathbf{X}_{pe}=[P_{i},D_{i},V_{i}]$ was passed through a fully connected (FC) layer, then concatenated along the channel dimension for fusion. Finally, another FC layer was applied for linear projection to generate the final representation:
$\mathbf{X}_{pe}=\mathbf{FC}\left(\operatorname{Concat}\left[\mathbf{FC}\left(P_{i}\right),\mathbf{FC}\left(D_{i}\right),\mathbf{FC}\left(V_{i}\right)\right]\right)\in\mathbb{R}^{T\times C_{d}}$ (3)
where: $C_{d}$ indicates the channel dimension of $X_{pe}$.
Ego-vehicle Motion Encoder. The motion of the ego-vehicle indicates a potential conflict with a pedestrian. Thus, speed $S_{i}=\left\{s_{t}\right\}\in\mathbb{R}^{T\times 1}$ and acceleration $Acc_{i}=\left\{acc_{t}\right\}\in\mathbb{R}^{T\times 1}$ of the ego-vehicle were utilized simultaneously to describe it. To represent the global $Acc_{i}$, it was calculated as:
$acc_{t}=\frac{\left(s_{T}-s_{0}\right)\times FPS}{3.6\times T}$ (4)
where: $FPS$ presents the frames per second depending on video attributes. Furthermore, similar to $\mathbf{X}_{pe}$, the $\mathbf{X}_{ev}$ was also encoded and fused by FC layers and concatenate operation:
$\mathbf{X}_{ev}=\mathbf{FC}\left(\text{ Concat }\left[\mathbf{FC}\left(S_{i}\right),\mathbf{FC}\left(Acc_{i}\right)\right]\right)\in\mathbb{R}^{T\times C_{d}}$ (5)
Skeleton Pose Encoder. The pedestrian skeleton pose reflects the action highly related to the crossing behavior, which can be represented as a graph $G=\{V,E\}$ naturally, where the joints are represented as nodes $V$ and limbs as edges $E$. $N=20$ keypoints $K_{i}=\left\{K_{i}^{1},K_{i}^{2},\ldots,K_{i}^{N}\right\}\in\mathbb{R}^{T\times N\times 3}$ was utilized, with each point is represented as $K_{i}^{j}=\left\{\left(x_{t}^{\prime},y_{t}^{\prime},s_{t}\right)\mid j\in 1,\ldots,N\right\}$. The adjacency matrix $\mathbf{A}$ includes the skeleton naturally connections and self-connected matrix. To avoid redundancy with position, each absolute coordinates of keypoint $(x_{t}^{k},y_{t}^{k})$ in the image was normalized with its corresponding top-left bounding box coordinates $(x_{t}^{btl},y_{t}^{btl})$ as: $\left(x_{t}^{\prime},y_{t}^{\prime}\right)=\left(x_{t}^{k}-x_{t}^{btl},y_{t}^{k}-y_{t}^{btl}\right)$, and $s_{t}$ is the confidence score. Inspired by the convolution in graph structure [21], four stacked GCN blocks with residual connections were proposed as feature extractors of $K_{i}$, as shown in Fig. 2 (b). Furthermore, learnable edge importance was also applied in the network training. The entire process can be denoted as:
$\begin{gathered}\widehat{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{AD}{}^{-\frac{1}{2}}\\ \mathbf{X}_{i}=\text{ BatchNorm }\left(K_{i}\right)\\ \mathbf{H}_{0}=\sigma\left((\mathbf{E}_{0}\odot\widehat{\mathbf{A}})\mathbf{X}_{i}\mathbf{W}_{0}\right)\\ \mathbf{H}_{l}=\sigma\left((\mathbf{E}_{l}\odot\widehat{\mathbf{A}})\mathbf{H}_{l-1}\mathbf{W}_{l}+\operatorname{Conv2D}\left(\mathbf{H}_{l-1};\mathbf{W}_{cl}\right)\right)\\ \end{gathered}$ (6)
where: $\mathbf{D}$ is the diagonal node degree matrix. $\mathbf{E}_{l}$ represents the learnable edge matrix that align with the dimension of $\mathbf{A}$ in layer $l$ ($l=1,2,3$), while $\mathbf{H}_{l}$ means the feature map. $\sigma(\cdot)$ indicates the ReLU activate function, and $\odot$ is Hadamard product operation. Further, the last layer $\mathbf{H}_{3}$ was flattened to $\mathbb{R}^{T\times(N\times d_{hid})}$ and linear projected as the output of $\mathbf{X}_{ke}$:
$\mathbf{X}_{ke}=\mathbf{FC}\left(\text{ Flatten }\left(\mathbf{H}_{3}\right)\right)\in\mathbb{R}^{T\times C_{d}}$ (7)
where: $d_{hid}$ is the dimension of hidden layers in GCN. After encoder layers, the $\mathbf{X}_{pe}$,$\mathbf{X}_{ev}$,and $\mathbf{X}_{ke}$ were concatenated along the channel dimension for feature fusion, with another FC layer mapping to the Transformer embedding space.
$\mathbf{X}=\mathbf{FC}\left(\operatorname{Concat}\left[\mathbf{X}_{pe},\mathbf{X}_{ev},\mathbf{X}_{ke}\right]\right)\in\mathbb{R}^{T\times C_{d}}$ (8)
So far, $\mathbf{X}$ comprises features from all modalities. To capture long-range dependencies, Transformer encoder layers [22] were employed by leveraging the self-attention mechanism to model the temporal dimension. Additionally, positional encoding (PE) was applied to preserve temporal order. After stacking several Transformer encoder layers, the output $\mathbf{X}_{\text{out}}$ was flattened and passed through an MLP head with a single fully connected layer for binary classification.
$\hat{Y}_{i}=\text{softmax}\left(\mathbf{FC}\left(\text{ Flatten }\left(\mathbf{X}_{\text{out }}\right)\right)\right)$ (9)
III Experiments
III-A Datasets and Implementation Details
The effectiveness of the proposed method was evaluated using two large-scale datasets in the context of autonomous driving – JAAD [23] and PIE [24]. Both datasets are annotated with the crossing point in time, enabling the TTE calculation. The video frame rate from the two datasets is 30 FPS. Instead of ego-vehicle speed, five ego-vehicle motion states were used in the $\mathbf{X}_{ev}$ of JAAD, as velocity data are not available. The data used in the PIE experiments was split into train, validation (val), and test at a ratio of 0.5:0.1:0.4, with random splitting based on person IDs. For comparison with existing methods (Table I and II), the performance following the data split in [3] was also evaluated, with Set 01, 02, 04 for training, Set 05, 06 for validation and Set 03 for testing. Each video set contains a continuous driving recording segmented into multiple 10-minute clips. For the JAAD dataset, the JAAD_all subset and adhered to the default data split were used [3]. Furthermore, 0.5 second ($T=16$) observations were utilized to predict crossing behavior after 1-2 seconds ($TTE\in[30,60]$), sampled data based on the sliding window with an overlap of 0.6, 0.8 for PIE and JAAD dataset, respectively.
In the network architecture, the dimension of three encoder layers $C_{d}$, hidden layers in GCN $d_{hid}$, and hidden size in Transformers are all set to 64. With the concatenate operation in the feature fusion block, the channel dimension of the fused feature is $1923\times C_{d}=$ 192. In the transformer encoder, $dropout=0.1$, 4 attention heads, and 4 stacked layers were applied to extract temporal information. The reference lines in PDM were set to one line between points: $A=(0,H)$ and $B=(W/2,Y_{min})$ while another between: $B=(W/2,Y_{min})$ and $C=(W,H)$, as shown in Fig. 2 (a), where $Y_{min}$ is the minimum $y_{t}$ of all samples in each dataset. The minor differences in coordinates have little impact on model performance in subsequent experiments, as shown in Table III. The keypoints were generated by the pre-trained Alphapose [25], while the additional neck, hip, and body center points were further calculated based on the average of adjacent points.
The cross-entropy loss [26] was utilized for network training. For the PIE dataset, we trained for 32 epochs with a batch size of 128, using the Adam optimizer with an initial learning rate of $1e^{-4}$, and employed the ReduceLROnPlateau scheduler to halve the learning rate if the validation loss do not decrease by $1e^{-4}$ over 8 epochs. In JAAD, the AdamW optimizer with a constant learning rate of $5e^{-5}$ was applied for 32 epochs training, with a batch size of 64. Additionally, a weight decay of $1e^{-4}$ was applied on both datasets to mitigate overfitting. All experiments were conducted on a server equipped with an NVIDIA TITAN X GPU with 12 GB of memory.
Figure 3: Visualizations of predicted pedestrian crossing intentions in challenging scenarios. Displayed are the crossing probabilities with and without our proposed PDM, where values above 50% indicate crossing. Our method performs robustly under complex conditions, including irregular intersections (a) and curved roads (e), while also capturing long-range pedestrian behavior (b), and ambiguous roadside behavior (c) and (d). Failure cases typically falling into the parallel crossing behavior during vehicle turning, keypoints outliers, and the pedestrians yielding behavior to the ego-vehicle, as shown on the right.
TABLE I: Performance comparisons of GTransPDM with state-of-the-art methods on the PIE and JAAD datasets. Best results in bold, while the second-best underlined.
Models PIE JAAD
[…]
BiPed[27] 0.91 0.90 0.85 0.82 0.83 0.79 0.60 0.52
Pedestrian Graph+ [28] 0.89 0.90 0.81 0.83 0.86 0.88 0.65 0.58
MultiRNN [29] 0.83 0.80 0.71 0.69 0.79 0.79 0.58 0.45
SFRNN [13] 0.82 0.79 0.69 0.67 0.84 0.84 0.65 0.54
SingleRNN [18] 0.81 0.75 0.64 0.67 0.78 0.75 0.54 0.44
PCPA[3] 0.87 0.86 0.77 / 0.85 0.86 0.68 /
TrouSPI-Net[30] 0.88 0.88 0.80 0.73 0.85 0.73 0.56 0.57
IntFormer[10] 0.89 0.92 0.81 / 0.86 0.78 0.62 /
FF-STP[4] 0.89 0.86 0.80 0.79 0.83 0.82 0.63 0.51
PIT [31] 0.91 0.92 0.82 0.84 0.87 0.89 0.67 0.58
GTransPDM (w/o $X_{ke}$) 0.92 0.90 0.86 0.85 0.84 0.72 0.54 0.56
GTransPDM 0.90 0.87 0.82 0.86 0.87 0.78 0.64 0.64
TABLE II: Model efficiency comparisons with recent models.
Model Acc (PIE) Acc (JAAD) Params (Millions) Inference Time (ms)
Pedestrian Graph+ [28] 0.89 0.86 0.07 5.47
PedGNN [32] 0.68 0.80 0.58
PCPA [3] 0.87 0.85 31.165 38.6
PIT [31] 0.91 0.87 21.2 4.8
GransPDM (w/o $X_{ke}$) 0.92 0.84 0.13 0.03
GransPDM 0.90 0.87 0.23 0.05
TABLE III: The impact of reference line coordinates.
$Y_{A},Y_{C}$ PIE JAAD
[…]
$H$ 91.21 88.13 81.61 87.43 78.16 64.00
$H-100$ 91.24 87.70 81.42 87.25 77.81 63.46
$H-200$ 91.16 87.65 81.29 87.36 77.74 63.52
TABLE IV: Ablation study on multi-modalities. Ours in bold. Left: PIE dataset, Right: JAAD dataset.
Encoders Choice
$X_{pe}$
$X_{ev}$
$X_{ke}$
[…]
F1$\uparrow$ 75.04/52.41 66.78/0.00 57.98/52.17 79.57/54.16 81.61/64.00
[…]
TABLE V: Impact on features used in $X_{pe}$. $P_{i}(C)$: coordinates variation in PDM, $P_{i}(R)$: area ratio in PDM, $P_{i}(\Delta d)$: real depth variation. Best in bold and second-best underlined in each dataset.
$X_{pe}$ Choice
$D_{i}$
$V_{i}$
$P_{i}(C)$
$P_{i}(R)$
$P_{i}(\Delta d)$
[…]
F1$\uparrow$ 78.80/54.22 77.35/54.11 80.33/55.19 81.61/64.00 81.93/54.65
[…]
III-B Evaluations
The effectiveness of our method was evaluated using Accuracy, Area Under Curve (AUC), F1 Score, Precision (P), and Recall (R), which are commonly used in existing studies. Table I and Table II compare our method and efficiency with recent approaches from the past five years, including Transformer-based IntFormer [10] and PIT [31], the graph-based Pedestrian Graph+[28], and several RNN-based methods. Except for Pedestrian Graph+ and BiPed, others show the same configurations as ours, following [3]. While PIT [31] captures traffic agent interactions using a ViT-based  [33] network with 16 stacked Transformer layers, incurring high computational costs, our lightweight network achieves superior performance, reaching 0.92 accuracy and 0.86 F1 in PIE using only position and ego-vehicle motion encoders, with 0.86 precision when incorporating the pose encoder. However, since [3] select samples following the sequence of long-period continuous videos, showcasing less randomness and performance degradation of $X_{ke}$ in our model. Thus, we conducted a 5-fold validation to further elucidate the contribution of $X_{ke}$ in Table VI of the Supplementary material and Table IV. Regarding JAAD, our method achieves an accuracy of 0.87 and a precision of 0.64, surpassing most existing approaches.
As an important component in the PDM, we evaluated the model’s performance with different reference line coordinates (see Table III). Since the values $(\Delta x_{t}-\Delta x_{t-1},\Delta y_{t}-\Delta y_{t-1})$ depend solely on the line’s slope, we varied the vertical coordinates of points $A$ and $C$ during testing. Our results indicate that slight variations in the line slope have minimal impact on performance, thereby underscoring the robustness in enhancing lateral pedestrian movement recognition. To evaluate the effectiveness of each network component, ablation studies were conducted on the modalities and features used in $X_{pe}$, as shown in Table IV and V. Left and right values represent metrics in PIE and JAAD respectively. In Table IV, the position encoder proves most effective in PIE, whereas skeletal information plays a critical role in JAAD. In contrast to the results in Table I, incorporating skeletal information in PIE leads to a 1-2% performance improvement. Overall, the model performs best when all encoders are combined. Table V further illustrates the different features used in $X_{pe}$. Our proposed PDM brings 4.26% and 9.89% F1 improvements in PIE and JAAD respectively. To compare it with real depth-based methods, we computed the metric depth variation $\Delta d=d_{t}-d_{t-1}$ using Depth Anything [34], a SOTA monocular depth estimation model, as a substitute for $R$ in PDM. While the model achieves comparable performance on the PIE dataset, it shows a significant drop on JAAD. This is mainly due to the limitations of monocular depth estimation, such as sensitivity to occlusion and failure in scenes with poor visibility. Fig. 3 presents challenging scenarios in PIE and JAAD datasets. Notably, in Fig. 3 (b), the ego-vehicle is traveling at high speed while the pedestrian remains at a distance, making it particularly challenging to detect the pedestrian’s approach. Owing to our PDM, the model reliably identifies crossing behavior, thereby ensuring a safe response by AVs. Beyond behavior prediction, the core idea of our model also facilitates human behavior understanding in mobile sensing scenarios, such as capturing fine-grained pose changes and providing a lightweight alternative to depth sensing.
IV Conclusion
In this study, we propose GTransPDM, a Graph-embedded Transformer with a Position Decoupling Module (PDM) for pedestrian crossing intention prediction. It effectively fuses multi-modal features with position, ego-vehicle motion, and pedestrian pose, ensuring a lightweight, real-time application. The PDM enhances pedestrian lateral movement recognition and captures depth cues, while a GCN-enhanced Transformer models spatio-temporal skeleton dynamics. Our approach predicts crossing intent 1–2 seconds in advance, achieving 92% accuracy on PIE and 87% on JAAD, with a runtime of 0.05ms, outperforming existing state-of-the-art methods.